OpenStudy (anonymous):

Use elimination to solve the following system of equations. 4x+7y=-16 -12x-7y=-8 A.(7, -8) B.(-4, 3) C.(8, -7) D.(3, -4)

OpenStudy (rajee_sam):

Elimination is you eliminate one variable by manipulation the equations. In the above set of equations if you just add them together variable y gets eliminated. because one has +7y and the other has -7y. Now if you add both equations you get 4x +( -12x) + 7y + (-7y) = -16 +( -8) -8x = -24 x = -24/-8 = 3 Now only answer choice is d (3, -4) pf you had 2 choices that had x = 3, then substitue x=3 in one equation and find y

OpenStudy (anonymous):

The y's cancel and your left with: 4x=-16 -12x=-8 Add down and you get: -8x=-24 Divide and you get x=3 plug in 3 into x in the beginning equation and the only possible choice is D